OpenStudy (anonymous):

Solve. 4x + 2 ≥ –6 A. x ≤ –2 B. x ≤ –1 C. x ≥ –2 D. x ≥ –1

OpenStudy (anonymous):

help @iPwnBunnies

hartnn (hartnn):

start by subtracting 2 from both sides, what do u get ?

OpenStudy (anonymous):

2 and 4

OpenStudy (anonymous):

some one help

OpenStudy (anonymous):

help @link2122

hartnn (hartnn):

4x+2-2 \(\ge \) -6-2 so, 4x \(\ge \) -8 got this ? now divide both sides by 4, what'd u get ?

OpenStudy (anonymous):

1 and 2

OpenStudy (anonymous):

1 and -2

OpenStudy (anonymous):

1 and -2

OpenStudy (anonymous):

whatevs

hartnn (hartnn):

yes, so u get \(x \ge -2\) and its correct :)

OpenStudy (anonymous):

thx

hartnn (hartnn):

welcome ^_^